Given y = 3/(5√x) . The expression for y2 is:

A. 6/(10x2)

B. 9/(25x2)

C. 6/(10x)

D. 9/(25x)


1
Expert's answer
2021-05-25T09:17:19-0400
"y^2=\big(\dfrac{3}{5\sqrt{x}}\big)^2=\dfrac{3^2}{5^2(\sqrt{x})^2}=\dfrac{9}{25x}, x>0"

D. 9/(25x)